Lemon drops worth $1.90 per pound are to be mixed with jelly beans that cost $1.20 per pound to make 100 pounds of a mixture worth $1.48 per pound. How many pounds of each candy should be used?
Found 2 solutions by checkley77, amalm06:
Answer by checkley77(12844)

1.9X+1.2(100-X)=100*1.48
1.9X+120-1.2X=148
.7X=148-120
.7X=28
X=28/.7
X=40 POUNDS OF LEMON DROPS ARE USED.
100=40=60 POUNDS OF JELLY BEANS ARE USED.
PROOF:
1.9*40+1.2*60=148
76+72=148
148=148

Answer by amalm06(224)

Using the method of alligation,
L=1.90, J=1.20
Then 1.90-1.48=0.42 and 1.48-1.20=0.28
So that L/J=0.28/0.42=2/3
Number of lemons=(100)(2/5)=40 (Answer)
Number of jelly beans=100-40=60 (Answer)